Q1lov a nov Hilr Tor 202 3204 sion b JOV TiH a] what is the current through the lo vemf near bottom left ? b] What is the poten
Answer #1

* I 20 + R 20 ox lov с # I IZ IL کمه Ans: In the ciscuit diagsamo two 20.2 desistances are connected in parallel Let R be theContinued 211 + I2 = 21 -It - 21% = -2 I = 2 A IL -I, = 0 I = I2 Now flom equation 0 - 21; +I=8 3 I = 2 = .667 A from the diaProblem is solved using the concept of parallel combination of resistance, KCL, KCL and ohm's law.